Circular No. 9081

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 T2 (LA SAGRA) R. Stoss, Visnjan, Croatia, reports the discovery of a diffuse comet

with no obvious tail on CCD images taken by J. Nomen in the course of the “La Sagra Sky Survey” with a 0.45-m f/2.8 reflector (discovery observation tabulated below). Thirty stacked 90-s follow-up CCD exposures with a 0.40-m f/10 reflector at La Sagra on Oct. 12.94–12.96 UT show a dropshaped coma elongated to the southwest with a 20′′ tail in p.a. 240^{o}; the coma is very condensed, almost asteroidal, and the tail is much fainter. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, R. Miles (Stourton Caundle, Dorset, England, 0.28-m f/9.4 Schmidt-Cassegrain reflector) writes that 127 stacked 30-s CCD frames taken on Oct. 13.0 show weak cometary appearance with a narrow tail 25′′ long in p.a. 230^{o}.

2009 UT α_{2000} δ_{2000} Mag. Observer
Oct. 12.87132 1^{h}44^{m}15.^{s}82 +21^{o}49′42′′.0 17.0 Nomen

The available astrometry (including pre-discovery Catalina astrometry from Sept. 18, 25, and Oct. 9 found by MPC staff), the following preliminary elliptical orbital elements, and an ephemeris appear on MPEC 2009-T39.

T = 2010 Jan. 12.815 TT  ω = 215.^{o}448
e = 0.76875  Ω = 215.987 }2000.0
q = 1.75499 AU  i = 28.102
a = 7.58906 AU n^{o} = 0.047144 P = 20.9 years

COMET P/2009 Q4 (BOATTINI) Additional astrometry have shown that this comet (cf. IAUC 9069) is

of short period, as indicated by the following elements from MPEC 2009-T18: T = 2009 Nov. 19.9438 TT, e = 0.578664, q = 1.320671 AU, ω = 320^{o}.0361, Ω = 127^{o}.6603, i = 10^{o}.9681 (equinox 2000.0), P = 5.55 years.

COMET P/2009 Q5 (McNAUGHT) Additional astrometry have shown that this comet (cf. IAUC 9070) is

of short period, as indicated by the following elements from MPEC 2009-T19: T = 2009 Sept. 8.5621 TT, e = 0.609033, q = 2.918346 AU, ω = 209^{o}.2727, Ω = 160^{o}.1369, i = 40^{o}.9012 (equinox 2000.0), P = 20.39 years.

2009 October 13  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green